# Learning in Structured Stackelberg Games

### v2

@@ 3.1 Leader’s objective
Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{i},f_{i})\}_{t=1}^{T}$ the leader’s regret is
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
where $\pi_{h^{*}}(\mathbf{z})=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have gotten if she could (1) optimally predict the type of follower given the context and (2) play an optimal strategy against that type of follower. In Section˜3 we present an algorithm that is optimal with respect to this benchmark.
@@
An important part of our analysis relies on the leader being able to compute the optimal strategy for any fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution. Previous work  [7, 27] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees. For completeness, we formally state this result in Section˜A.1.
@@ 3.2 Littlestone Dimension
In this section, we highlight the concepts and definitions from multiclass online learning [22] that are relevant to our work. We start with the notion of shattering, which captures the following intuition: the richer a set of functions is, the more complex patterns it should be able to fit.
Definition 3.3 (Multiclass $\pazocal{H}$-Shattered Littlestone Tree).
@@
The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^3] We note that for binary labels, the Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes at the same depth of a shattered tree are the same. This definition is due to Daniely et al. [22] and it generalizes the classic definition of the Littlestone dimension for binary classification [30] to the multiclass setting.
Definition 3.4 (Multiclass Littlestone Dimension).
The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted as $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
Examples.
The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [30]. There are several examples of hypothesis classes that have finite Littlestone dimension and can, therefore, be (online) learned. These include the class of conjunctions, the class of decision lists, the class of linear separators with a margin (cf. Proposition 17 in [1]), and in general anything learnable in the mistake bound model.
In Daniely et al. [22], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [20]. However, it is a priori unclear whether learning to optimally predict the follower’s type is also necessary. In  Theorem˜3.5 we show that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information. The full proof can be found in Section˜A.2.
@@
There exists a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$, such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
@@
Our proof captures the following intuition: in a structured Stackelberg game, the follower types can be completely distinct from the perspective of the leader (i.e., any one irreducible to another), and the hypothesis class can be arbitrarily hard to learn. Yet, to truly know how difficult the leader’s learning task is, one has to account for how the parameters of the game interact with the hypothesis class. Our construction uses distinct follower types and the hypothesis class indicates that there is a true threshold that splits the context domain in two parts, each of which is predictive of a different follower type. Learning that true threshold can be arbitrarily hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), yet unnecessary if the two follower types happen to induce the same optimal strategy for the leader near the threshold.
This result highlights that the Littlestone dimension is completely blind to how the hypothesis class interacts with the utilities of the players, which motivates a new measure of complexity.
3.3 Stackelberg-Littlestone Dimension
In this section we provide an instance optimal learner for structured Stackelberg games. We start with the definition of node-weighted shattered trees.

### v3

@@ 3 Online learning
Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{t},f_{t})\}_{t=1}^{T}$ the leader’s regret is $R(T):=\sum_{t=1}^{T}r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$ where $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have obtained if she could correctly predict the follower’s type given the context and play optimally against it.
@@
Our analysis relies on the leader being able to compute the optimal strategy to play for any fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader may not be able to get an exact solution. Previous work [balcan2015commitment, harrisregret] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of exposition, we assume that the leader can directly optimize over the entire simplex, without affecting her regret guarantees. A formal statement of this result appears in Section˜B.1.
@@ 3.1 Littlestone dimension
Given the structure induced by the finite set of follower types, a natural idea is to directly predict the follower’s type from the context, and play the corresponding optimal strategy. This suggests using tools from online (multiclass) classification to characterize the statistical complexity of the problem. In this section, we will show that this approach comes up short.
@@
We begin by recalling several concepts from online learning that are relevant to our work, starting with the notion of shattering, which captures the intuition that richer hypotheses classes can encode more complex label patterns.
Definition 3.3 (Multiclass Littlestone Tree).
A Littlestone tree is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in a finite set $[K]$. We say that the tree is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path that traverses nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{i}$ is $h(\mathbf{z}_{i})$ for all $i$.
Example 3.4.
Consider the class of permutations
$\mathcal{H}_{3}=\left\{\,h:\{1,2,3\}\to\{f^{(1)},f^{(2)},f^{(3)}\}\ \middle|\ h\text{ is a bijection}\,\right\}.$
Figure˜1 shows an example of a Littlestone Tree that is shattered by this class.
Figure 1: $\mathcal{H}_{3}$-shattered Littlestone Tree showing a consistent hypothesis for each root-to-leaf path.
The multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^4] For binary labels and when all nodes at the same depth of a shattered tree are the same, the Littlestone dimension recovers the definition of the well-studied VC-dimension. This definition is due to pmlr-v19-daniely11a and it generalizes the classic definition of the Littlestone dimension for binary classification [littlestone1988learning].
Definition 3.5 (Multiclass Littlestone Dimension).
The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$ such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
The finiteness of the Littlestone dimension of a hypothesis class is necessary for a hypothesis class to be online learnable [littlestone1988learning]. Some examples of classes with finite Littlestone dimension are conjunctions, decision lists, linear separators with a margin, and in general anything learnable in the mistake bound model (cf. balcannotes:mistakebound, Proposition 17 in Alon21:partial).
pmlr-v19-daniely11a show that the multiclass Littlestone dimension of a hypothesis class characterizes learnability in online multiclass classification, that is, it determines the number of mistakes an optimal online learner will make in the worst-case. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can compute the strategy that maximizes her utility by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [conitzer2006computing].
However, it is a priori unclear whether learning to optimally predict the follower’s type is not only sufficient but also necessary. We now show that this is not the case, by constructing a structured Stackelberg game instance with an infinite Littlestone dimension that can be solved without learning.
Theorem 3.6.
There exists a Stackelberg game $\pazocal{G}$ and a hypothesis class $\pazocal{H}$ such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
Proof sketch.
The key idea is that, although the follower types can be distinct from the leader’s perspective (Remark˜2.1) and the hypothesis class can encode arbitrarily complex type-prediction problems, the leader’s downstream utility optimization problem may be much simpler. Our construction uses two follower types and a hypothesis class, $\pazocal{H}$, consisting of linear threshold functions. Learning the true threshold can be hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), but is unnecessary: near the threshold the two follower types induce the same optimal strategy for the leader.  Figure˜2 provides an illustration of the Stackelberg game we use in the proof. ∎
Figure 2: Follower types as a function of the context in the construction of Theorem˜3.6. While $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the leader’s optimal strategy only depends on whether $\mathbf{z}[2]<0.5$.
This result shows that a finite Littlestone dimension is not necessary for online learning in our setting, as it only measures the complexity of predicting the follower types, not how those predictions affect the leader’s utilities.
3.2 Stackelberg Littlestone dimension
We now define the Stackeleberg Littlestone dimension, which we prove characterizes online learnability in structured Stackeleberg games.
In the definitions that follow we consider trees with nodes labeled by elements in $\pazocal{Z}$ and edges corresponding to different labels in $[K]$. A depth-$d$ tree, $\mathcal{T}_{d}=\{\mathbf{z}_{s}:s\in S_{d}\}$, is indexed by sequences $S_{d}\subseteq\{\varnothing\}\cup\{s\in[K]^{\ell}:1\leq\ell\leq d\}$, where $\mathbf{z}_{\varnothing}$ denotes the root node.
Definition 3.7 (Stackelberg Littlestone (SL) Tree).
A SL tree, $\pazocal{T}_{d}$, is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in $[K]$. Each node $\mathbf{z}_{s}\in\mathcal{T}_{d}$ has a weight $\rho_{s}$, where $\rho_{s}=0$ if $\mathbf{z}_{s}$ is a leaf node and $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]:sj\in S_{d}}\left(r(\mathbf{z}_{s},\mathbf{x},f^{(j)}))+\rho_{sj}\right)$ otherwise. We say that the tree is shattered by a hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$ if for every root-to-leaf path that traverses nodes $\mathbf{z}_{\varnothing},\dots\mathbf{z}_{s\leq d}$, there exists $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{s_{\leq i}}$ is $h(\mathbf{z}_{s_{\leq i}})$ for every $i$.
Example 3.8.
Consider the hypothesis class of permutations
$\mathcal{H}_{3}=\left\{\,h:\{1,2,3\}\to\{f^{(1)},f^{(2)},f^{(3)}\}\ \middle|\ h\text{ is a bijection}\,\right\}$
and a Stackelberg game with 3 actions for both players. The leader’s utility matrix is $U_{L}=\mathbb{I}_{3\times 3}$ and follower $f^{(i)}$’s utility matrix is $U_{f^{(i)}}=\mathbf{1}e_{i}^{\top}$, i.e. she receives utility 1 if and only if she plays action iiFigure˜3 shows a corresponding SL tree.
Figure 3: $\mathcal{H}_{3}$-shattered SL-Tree showing the consistent hypothesis for each root-to-leaf path. Leaf nodes have weight 0, nodes at depth 11 have weight $1/2$ and the root node has weight $1/2+2/3$.
Definition 3.9 (Stackelberg Littlestone (SL) dimension).
The SL dimension of a hypothesis class $\mathcal{H}$ for Stackelberg game $\pazocal{G}$ is defined as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H}):=\sup\{\kappa\geq 0:$ there exists an SL tree shattered by $\pazocal{H}$ under $\pazocal{G}$ with root node weight $\kappa\}$.
Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value. If for every $\kappa\geq 0$ there exists an SL tree with root node weight at least $\kappa$, then $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=\infty$.
We now show that no deterministic algorithm can achieve regret lower than the SL dimension.
Theorem 3.10.
For any hypothesis class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$, there exists an adversarial sequence of contexts and follower types such that any deterministic algorithm suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.
Proof sketch.
The key idea in the proof is that the weight assigned to each node of an SL tree measures how much regret the adversary can still guarantee from that point onward. At the root, this weight captures the minimum regret the learner must suffer no matter which strategy she plays. If the learner chooses the strategy that minimizes immediate regret, the adversary selects the follower type that leads into a subtree with large remaining weight, ensuring future regret. If the learner deviates to reduce future regret, the adversary instead extracts more regret immediately.
By repeating this reasoning along a root-to-leaf path, the adversary can force the learner to incur regret equal to the root weight of the tree. Since the SL dimension is defined as the supremum root weight over all such shattered trees, no deterministic algorithm can achieve regret smaller than the SL dimension (up to an arbitrarily small additive term). ∎
We now present an instance-optimal algorithm that achieves regret equal to the SL dimension. We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), as it shares some similarities with the Standard Optimal Algorithm (SOA) from online multiclass prediction [daniely2014:optimal]. In each round, SSOA tracks the subset of hypotheses that are consistent with the history so far. The algorithm then plays the mixed strategy that minimizes the worst-case instantaneous regret, plus the SL dimension of the induced hypothesis class. The latter captures the difficulty of the learning task that remains after the current round ends.
1:Input: Stackelberg game $\pazocal{G}$, hypothesis class $\pazocal{H}$
2: Initialize version space $V_{1}=\pazocal{H}$
3:for $t\in[T]$ do
4:  Observe context $\mathbf{z}_{t}$
5:  Compute optimal utility for each type $i\in V_{t}(\mathbf{z}_{t})$ $u_{*}^{(i)}=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))\;$
@@ 3.2 Stackelberg Littlestone dimension
7:  Observe follower type $f_{t}$
8:  Update version space $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
9:end for
ALGORITHM 1 SSOA
@@
For any Stackelberg game $\pazocal{G}$ and every hypothesis class $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
@@
Our proof follows a high-level idea similar to the analysis of SOA in online prediction: With each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases. Similarly, we show that when the learner incurs instantaneous regret, the SL dimension of the induced hypothesis class decreases by at least this amount.∎
3.3 Relationship with the Littlestone dimension
We have so far shown that the SL dimension characterizes learnability in structured Stackelberg games. We now examine how the it relates to online multiclass prediction.
Lemma 3.12.
For any hypothesis class $\pazocal{H}$ and any Stackelberg game $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
To make the relationship between the two dimensions more concrete, consider the following example, which is a generalization of Examples 3.4 and 3.8 to $n$ dimensions.
Example 3.13.
Consider a setting with $n$ contexts, $n$ follower types, and the class of permutations $\pazocal{H}_{n}$. Note that $\textnormal{Ldim}(\pazocal{H}_{n})=n-1$, since an adversary can force a mistake in every round that she can query. A tree of depth $n$ cannot be shattered because once the labels of the first n1n-1 contexts have been observed, the label of the last context is fixed.
Now consider a Stackelberg game where the leader’s utility matrix is $U=\mathbf{I}_{n\times n}$ and follower type $i$’s utility matrix is $U_{i}=\mathbf{1}_{n}e_{i}^{\top}$. We will show that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H}_{n})=n-H_{n}$, where $H_{n}$ is the $n$-th harmonic number.
Consider the SL tree that witnesses the SL dimension. Due to symmetry, each node at depth $d$ will have ndn-d outgoing edges, one for each follower type not encountered in the path so far. Each node at depth n2n-2 will have weight $1/2$, since the optimal utility of the leader against any follower type is $1$, the optimal utility against any follower chosen adversarially from a set of two types is $1/2$, and any deeper nodes must have only one outgoing edge (and therefore weight 0). More generally, for a node $s$ at depth $d$ we have
$\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{j_{1},\dots,j_{n-d}\}}\biggl[r(\mathbf{z},\mathbf{x},f^{(j)})+\sum_{j=2}^{n-d-1}1-\frac{1}{j}\biggr],$
which simplifies to $n-d-\sum_{j=1}^{n-d}\frac{1}{j}$.
Note that while any deterministic learner will make at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning $\pazocal{H}$, mispredicting the follower’s type need not lead to suboptimal utility for the leader. Therefore, Lemma˜3.12 does not immediately imply that the optimal multiclass prediction algorithm (SOA) would incur suboptimal regret in our setting. Indeed, in Theorem˜3.6 we constructed a game instance where the leader can mispredict the follower’s type at every timestep, yet still achieve optimal utility. We conclude this section by showing that there are indeed game instances where SOA will suffer larger regret than SSOA.
Theorem 3.14.
There exists a hypothesis class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$ such that running SOA on $\pazocal{H}$ results in suboptimal utility for the leader.
Proof sketch.
The game construction exploits the mismatch between the label prediction objective of SOA and optimizing for leader utility. Specifically, the game is designed such that SOA focuses on parts of the hypothesis class that are combinatorially complex (large Littlestone dimension) but strategically irrelevant. As a result, SOA unnecessarily suffers utility loss over multiple rounds. ∎
4 Distributional learning
We now shift our focus from online learning to distributional learning. In this setting, contexts are drawn i.i.d. from an unknown distribution and as in the online setting, we assume that there exists a function $h^{*}\in\pazocal{H}$ that correctly predicts each follower type given the context. The leader’s goal is to choose a strategy that will maximize her expected utility on a fresh sample drawn from the same distribution. A learning algorithm is a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$.
We consider the following notion of sample complexity, adapted to our Stackelberg setting.
Definition 4.1 (PAC cut-off sample complexity).
The PAC cut-off sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of function class $\pazocal{H}$ w.r.t. Stackelberg game $\pazocal{G}$ is $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the smallest integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathcal{D}$ on $\pazocal{Z}$, and any target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$, $\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(\mathbf{z};S),f^{(h^{*}(\mathbf{z}))})>\gamma]$, is at most $\epsilon$ with probability at least $1-\delta$.
While we present our results with respect to the PAC cut-off sample complexity, any bound with respect to this benchmark immediately implies a bound on the familiar PAC sample complexity [valiant1984theory] as well.  [^5] See Lemma 1 in Attias23:regression.
4.1 Sample complexity lower bound
We now introduce the $\gamma$-valued Stackelberg-Natarajan dimension (henceforth SN dimension), which we show characterizes the sample complexity lower bound for distributional learning in structured Stackelberg games. Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting. The SN dimension is inspired by the Natarajan dimension from multiclass prediction [natarajan1989learning, daniely2014:optimal] and the $\gamma$-Natarajan dimension from realizable regression [Attias23:regression].
We start by defining what it means for a set of contexts to be $\gamma$-SN-shattered. Intuitively, a set is $\gamma$-SN-shattered if the hypothesis class $\pazocal{H}$ can fit complex patterns of follower types on the set. Furthermore, these follower types must be sufficiently different: no leader strategy can incur loss less than $\gamma$ if the type is chosen adversarially among them.
Definition 4.2 ($\gamma$-SN-shattered set).
A set $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}\subseteq\pazocal{Z}$ is $\gamma$-SN-shattered by hypothesis class $\pazocal{H}$ with respect to Stackelberg game $\pazocal{G}$ if there exist functions $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
  1. 1.
    $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma$ for every $\mathbf{z}_{i}\in S$, and
  2. 2.
    for all $b\in\{0,1\}^{n}$, there exists an $h_{b}\in\pazocal{H}$ such that
    $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1.\end{cases}$
Definition 4.3 ($\gamma$-valued SN dimension).
The $\gamma$-valued SN dimension, $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
We emphasize that that the two functions, $g_{0}$ and $g_{1}$ in Definition˜4.2 must be $\gamma$-far from each other in a minmax sense: at each context the leader cannot choose a mixed strategy that performs well against both $g_{0}(\mathbf{z}_{0})$ and $g_{1}(\mathbf{z}_{0})$. This enables a distribution-independent lower bound that holds for every learning algorithm. We prove our lower bound in Theorem˜4.4 by adapting the argument of Attias23:regression to our setting.
Theorem 4.4.
Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)$ such that $\delta<\epsilon$, then
Proof sketch.
The high-level idea is to reduce distributional learning in structured Stackelberg games to a hard identification problem over a $\gamma$-SN-shattered set. Such a set contains many contexts where the hypothesis class can realize two different follower types such that no leader strategy can simultaneously perform well against both.
The distribution we consider on this shattered set has most probability mass on one “easy” context and small but non-negligible mass spread over many “hard” contexts. With limited samples, the learner is unlikely to observe enough of the hard contexts to correctly distinguish the follower type. Since the hypothesis class can label these unseen contexts arbitrarily while remaining consistent with the data, the learner’s strategy on them is effectively a guess.
Because each unseen context independently forces a $\gamma$-loss with constant probability, the learner’s expected cut-off error remains large unless the number of samples scales linearly with the size of the $\gamma$-SN-shattered set. Standard concentration arguments then show that achieving small error with high probability requires a sample size proportional to the $\gamma$-SN dimension. ∎
4.2 Sample complexity upper bound
We conclude our treatment of the distributional setting by introducing the $\gamma$-valued Stackelberg-Graph dimension (henceforth SG dimension), which provides a sufficient condition for distributional learning in structured Stackelberg games. The SG dimension is a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [natarajan1989learning, daniely2014:optimal] and regression [Attias23:regression] to our setting.
Definition 4.5 ($\gamma$-SG-shattered set).
A set $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}\subseteq\pazocal{Z}$ is $\gamma$-SG-shattered by hypothesis class $\pazocal{H}$ with respect to Stackelberg game $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for every $b\in\{0,1\}^{n}$, there exists an $h_{b}\in\pazocal{H}$ satisfying
  1. 1.
    $h_{b}(\mathbf{z}_{i})=g(\mathbf{z}_{i})\;\;\forall i\in[n]$ such that $b_{i}=0$, and
  2. 2.
    $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma$ for all $i\in[n]$ such that $b_{i}=1$.
Definition 4.6 ($\gamma$-valued SG dimension).
The $\gamma$-valued SG dimension, $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SG-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
We now present the main result of this section. Given a training sample $S$ of context–follower-type pairs, Algorithm˜2 first computes the subset of hypotheses that are perfectly consistent with $S$. Then, on a new context, the algorithm plays the leader’s strategy that is optimal against an adversarially chosen follower type predicted at that context by any hypothesis consistent with $S$.
1:Input: Stackelberg game $\pazocal{G}$, hypothesis class $\pazocal{H}$, sample $S=\{(\mathbf{z}_{i},f_{i})\}_{i\in[n]}$, and test context $\mathbf{z}\in\pazocal{Z}$
2: Compute $\pazocal{H}|_{S}=\{h\in\pazocal{H}:h(\mathbf{z}_{i})=f_{i}\;\forall i\in[n]\}$
3: Compute $F=\pazocal{H}|_{S}(\mathbf{z})=\{h(\mathbf{z})\in[K]:h\in\pazocal{H}|_{S}\}$
4: Play $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
ALGORITHM 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
We now show that the $\gamma$-SG dimension upper bounds the PAC sample complexity in our setting. The proof of Theorem˜4.7 follows from adapting that of Attias23:regression to our setting.
Theorem 4.7.
For any class $\pazocal{H}$, game, $\pazocal{G}$, and $\epsilon,\delta,\gamma\in(0,1)$, there is some constant $C_{1}$ such that Algorithm˜2 achieves
Proof sketch.
The key observation in our analysis is that the leader only needs to worry about contexts where different hypotheses would force meaningfully different leader utilities. Algorithm˜2 keeps all hypotheses that are consistent with the data and, given a new context, plays the strategy that is most robust to all follower types predicted by those hypotheses. If the learner performs poorly on a noticeable fraction of contexts, then there must exist many contexts where some hypothesis in the class disagrees with the true one and where this disagreement induces loss at least $\gamma$. Such contexts form a $\gamma$-SG-shattered set.
The definition of the $\gamma$-SG dimension limits the size of these hard sets. Using standard uniform convergence and symmetrization arguments, we show that once the sample size exceeds this dimension, no hypothesis consistent with the data can be $\gamma$-far from the true one on more than an $\epsilon$ fraction of contexts. As a result, the chosen strategy achieves near-optimal expected utility with high probability.∎
5 Conclusion and future directions
We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type. Our main results consist of a complete characterization of instance-optimal regret bounds in the online setting, and PAC sample complexity guarantees in the distributional setting. There are several exciting directions for future work.
Our main focus in this work is on the statistical complexity of learning in structured Stackelberg games. An interesting avenue for future research is to consider the computational complexity of learning in this setting. We conduct a preliminary exploration of this direction in in Appendix˜D, and find that our results open the door for efficient (albeit so far sub-optimal) learning algorithms in some settings.
We focus on the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor. The agnostic learning setting is a natural relaxation, where this assumption is not made and the goal is instead to do as well as the best hypothesis in the class. As such, providing a complete characterization for the agnostic case remains an important future direction.
6 Acknowledgments
K.F. is grateful to Brian Hu Zhang for helpful discussions on an earlier version of this paper.
Appendix A Applications beyond Stackelberg games
Our results also apply to the problems of learning to bid in simultaneous second-price auctions with side information and Bayesian persuasion with both public and private states. These settings are formally introduced in balcan2025nearlyoptimalbanditlearningstackelberg; we give an overview here. They are mathematically equivalent to the setting we study in the main body, with the exception of the form that the learner’s strategy space takes (detailed below).
A.1 Simultaneous second-price auctions with side information
This application is an extension of a setting in daskalakis2016learning to incorporate side information. In each round (or historical data point in the distributional setting), bidders simultaneously bid on a set of $m$ items. They receive the bundle of items for which they are the highest bidder, and they pay a price corresponding to the second-highest bid for each item.
The learner in this setting is a bidder whose strategy space consists of a set of possible bids for each item. The bidder’s valuation depends on the entire bundle of items they receive, e.g. the bidder may have a low valuation for item A and item B in silos, but a high valuation for the combination of the two. In addition to the bundle of items, the bidder’s valuation also depends on contextual information, e.g. different seasons or fashion trends may influence a bidder’s valuation for a bundle of clothing items.
Other players’ bids roughly correspond to different follower types, and we posit a hypothesis class mapping contextual information to different sets of other players’ bids. Our results carry over to this application as-is when the space of the other players’ bids is discrete; the only difference is that the leader’s strategy space is now the bid space $0,1]^{m}$ instead of a probability simplex.
A.2 Bayesian persuasion with both public and private states
In Bayesian persuasion [kamenica2011bayesian], a sender (analogous to the leader in Stackelberg games) strategically reveals payoff-relevant information about an underlying private state of the world to a receiver (analogous to the follower), who then takes a payoff-relevant action. This sub-section considers a generalization of Bayesian persuasion where there is some additional payoff-relevant information about the state which is common knowledge to both players. Receiver types map to follower types, and we posit a relationship between the common knowledge and the receiver’s type. As is the case in the previous sub-section, our results carry over here as-is, with the sender optimizing over a convex polytope strategy space instead of a probability simplex.
Appendix B Details and proofs from Section˜3
See 3.6
Proof.
We construct such a game explicitly. Let the context space be $\pazocal{Z}=\{0,1\}\times[0,1]$ and let the set of follower types be $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$. Consider the follower types $f^{(1)}$ and $f^{(2)}$ with utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ chosen such that for every context $\mathbf{z}\in\pazocal{Z}$, the leader’s optimal action differs, i.e. $\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$. Concrete payoff matrices for all follower types are shown in Table˜1 (the leader is the row player and the follower is the column player).
$F_{1}$ $F_{2}$
$L_{1}$ $1,0.5$ $0.5,0$
$L_{2}$ $0.25,1$ $0.75,0$
(a) Follower type $f^{(1)}$
$F_{1}$ $F_{2}$
$L_{1}$ $1,0$ $0.5,0.5$
$L_{2}$ $0.25,0$ $0.75,1$
(b) Follower type $f^{(2)}$
$F_{1}$ $F_{2}$
$L_{1}$ $1,1-\mathbf{z}[1]$ $0.5,0.5$
$L_{2}$ $0.25,0.5$ $0.75,\mathbf{z}[1]$
(c) Follower type $f^{(3)}$
$F_{1}$ $F_{2}$
$L_{1}$ $1,\mathbf{z}[1]$ $0.5,0.5$
$L_{2}$ $0.25,0.5$ $0.75,1-\mathbf{z}[1]$
(d) Follower type $f^{(4)}$
Table 1: Utility tables for Theorem˜3.6
We now construct the hypothesis class $\pazocal{H}=\{h_{\gamma_{1},\gamma_{2}}:\forall\;\;0<\gamma_{1}<0.5,\;0.5\leq\gamma_{2}<1\}$, where